OpenStudy (mayaal):

giving a medal and a fan!!!! Write the sum using summation notation, assuming the suggested pattern continues.

OpenStudy (mayaal):

options for answers are attached

OpenStudy (dangerousjesse):

So tiny..

OpenStudy (mayaal):

what do u mean?u can't see the options?

OpenStudy (dangerousjesse):

No, I can see them haha. I think it's funny how small they are.

OpenStudy (mayaal):

ok,can u try to open them in the new tab and then zoom in?

OpenStudy (dangerousjesse):

Yeah, I went ahead and opened them in paint. It's fine :)

OpenStudy (mayaal):

ok:)

OpenStudy (mayaal):

@ganeshie8 @amistre64 @satellite73

OpenStudy (mayaal):

@DangerousJesse do u know how 2 do this one?

OpenStudy (mayaal):

@amoodarya

OpenStudy (amoodarya):

yes but what is the question ?

OpenStudy (dangerousjesse):

I just don't understand this, namely because \[\infty \] is a concept, not a numerical value..

OpenStudy (mayaal):

the question is: Write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 + ... + 67

OpenStudy (dangerousjesse):

Oh that actually makes sense

OpenStudy (mayaal):

options for answers are attached

OpenStudy (dangerousjesse):

You're adding five every time.

OpenStudy (mayaal):

i forgot to copy the question.sorry guys!

OpenStudy (amoodarya):

a1=-8 a2=-3 a3=+2 ... so they are arithmetic sequence

OpenStudy (amoodarya):

and an=a1+(n-a)d so a2-a1=5 so an=-8+5n an=67 -8+5n=67 ---> n=15 so sigma (0 to 15) (-8+5n)

OpenStudy (mayaal):

how did u find out that an=-8+5n

OpenStudy (amoodarya):